OpenStudy (anonymous):

in function form 9y-4x-9=0

OpenStudy (anonymous):

subtract 9y from both sides and then divide both sided by -9

OpenStudy (precal):

9y=4x+9 y=(4/9)x + 1 f(x)=(4/9)x+1

OpenStudy (anonymous):

i think its...y=1=4/9x

OpenStudy (anonymous):

or y=f(x) = (4/9) x +1